# Stepwise Regression and Pre-trained Edge for Robust Stereo Matching

### v3

@@ V Conclusion
In this paper, we propose SR-Stereo, a novel stereo matching architecture. Our approach addresses the issue of distribution discrepancies across different datasets by predicting disparity clips. To enhance the accuracy of these disparity clips, we introduce a loss weight that is determined by the regression objective scale. Through extensive evaluations on datasets including SceneFlow, KITTI, Middlebury 2014, and ETH3D, we demonstrate that SR-Stereo achieves competitive performance in disparity estimation and state-of-the-art cross-domain generalisation performance. Moreover, we demonstrate that this stepwise regression architecture can be generalised to existing iteration-based methods to improve performance without changing the network structure.
Additionally, we propose a method called Domain Adaptation Based on Pre-trained Edge (DAPE) to address the issue of edge blurring in fine-tuned models with sparse ground truth. Specifically, we use the RGB image and the disparity predicted by the pre-trained stereo matching model to estimate the edge map of the target domain image. This eliminates the need for a complex edge estimation network. The generated edge map is then filtered to generate edge map background pseudo-labels. These pseudo-labels, along with the sparse ground truth disparity on the target domain, are used as supervision to jointly fine-tune the pre-trained stereo matching model. Experimental results at KITTI, Middbury 2014, and ETH3D show that DAPE significantly improves the disparity estimation performance of the fine-tuned model, especially in untextured and detailed regions.
In our future work, we will delve deeper into advancing stepwise regression model to achieve more powerful cross-domain generalisation performance. In addition, we plan to design a more advanced edge estimator and develop an improved method for filtering edge pseudo-labels. These enhancements will further enhance the domain adaptation performance of the fine-tuned model, enabling it to better handle challenging scenarios and improve disparity estimation performance.

### v4

@@ V Conclusion
In this paper, we propose a novel stereo method, SR-Stereo, and a novel fine-tuning framework, DAPE. SR-Stereo overcomes the domain discrepancies by predicting multiple range-controlled disparity clips to achieve better generalization performance and domain adaptation performance. Furthermore, for domain adaptation on sparse ground truth, the proposed DAPE uses generated edge pseudo-labels to provide additional supervision to the fine-tuned model. We have conducted extensive experiments on Sceneflow, KITTI, Middlebury, and ETH3D, and effectively demonstrated the advancements of SR-Stereo and the effectiveness of DAPE. Compared to the baseline, the proposed SR-Stereo achieves both better cross-domain performance and in-domain performance. This result is in complete contrast to many generalized stereo methods, which tend to have poor in-domain performance. In the future, we plan to implement a stereo method based entirely on stepwise regression by designing more components related to stepwise regression architecture.